OpenStudy (anonymous):

Factor completely: x^2+9x+18 I know my answer will be written in the form (x +or- ...) (x +or- ...) and i know i have to find what multiplies to be 18 and adds to be 9, but this is where I get stuck. I can't for the life of me think of what the heck would multiply to be 18 and add to be 9. 3 and 4, 9 and 2, and 18 and 1 wont work and I cant use negative numbers since it's all positive right? am i missing it completely? Is it super obvious or does it even have one i can use? Is it no solution? argh...

OpenStudy (anonymous):

Factors of 18: 9*2 = 18 6*3 =18 18*1 = 18 Now add them and see which are equal to 9. 9+2= 6+3= 18+1=

OpenStudy (anonymous):

And remember that, \[\huge 3\times4\neq 18\]

OpenStudy (anonymous):

oh wow i ..... completely forgot 6x3... good going me wow that was dumb

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thank you lmao

OpenStudy (anonymous):

No problem!